OpenStudy (pottersheep):

What are the smalelst values of x that would give whole numbers as a result? 13x+1/3

OpenStudy (pottersheep):

actually (13x+1)/3

OpenStudy (amistre64):

define small

OpenStudy (pottersheep):

Well, the answers are 2 . 5 and 8

OpenStudy (anonymous):

so you want \[\frac{13x+1}{3}\] to be a whole number is that correct?

OpenStudy (pottersheep):

2, 5

OpenStudy (pottersheep):

Yes please

OpenStudy (anonymous):

in other words \[13x+1\] must be a multiple of 3 right? try x = 2

OpenStudy (amistre64):

i spose whole number is a postive integer

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

Yep positive

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

Is there any ,ethod of doing it though? Except for trial and error?

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

*method

OpenStudy (anonymous):

for x = 2 you get \[13\times 2+1=27\]

OpenStudy (anonymous):

well once you have x = 2 the rest should be clear. and x = 2 was not too hard to find

OpenStudy (amistre64):

13x+1 = 1,3,6,9,12, etc

OpenStudy (amistre64):

err... not 1 but i menat 0 if im thinking right

OpenStudy (anonymous):

if x = 2 works that means \[13\times 2+1\] is divisible by 3, in fact it is \[3\times 9\] therefore since 3 and 13 are relatively prime the next one what will work is \[2+3=5\] try it

OpenStudy (amistre64):

\[\frac{13x+1}{3}=N\] \[13x+1=3N\] \[13x=3N-1\] \[x=\frac{3N-1}{13}\] where x is any whole number

OpenStudy (amistre64):

ugh ... where N is any whole number

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

hmmm

OpenStudy (pottersheep):

I need the 3 smallest numbers

OpenStudy (pottersheep):

Thanks both of you !!

OpenStudy (amistre64):

yep